OpenStudy (anonymous):

Write a formula for the surface area of a sphere in terms of C, the circumference of a great circle on the sphere.

OpenStudy (anonymous):

4 pi r ^2 = C circumference = 2 pi r = (4 pi r^2) /(2r)=C/2r

OpenStudy (anonymous):

see, that's what I thought, but the book came out with SA=C^2/pi

OpenStudy (anonymous):

let's wait and see if someone has another ans :D

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

C = 2 pi r -> C^2 = 4 pi^2 r^2 =>SA = 4 pi^2 r^2 / pi = 4 pi r^2

OpenStudy (anonymous):

which is SA = C^2/ pi

OpenStudy (anonymous):

how did you know to write C^2 = 4pi^2r^2?

OpenStudy (anonymous):

C = 2 pi r ->c^2 = (2pi r)^2 = 4 pi^2 r^2

OpenStudy (anonymous):